Sheila Williams, a medical secretary, earns $3,575 monthly for a 35-hour week. For overtime work, she receives extra pay at the regular hourly rate up to 40 hours and time and one-half beyond 40 hours in any week. During one semimonthly pay period, Williams worked 10 hours overtime. Only 2 hours of this overtime were beyond 40 hours in any one week. Compute the following amounts.

Round all divisions to two decimal places and use the rounded amounts in subsequent computations. Round your final answers to the nearest cent.

a. The regular semimonthly earnings $
b. The overtime earnings $
c. The total earnings $
Answer #1

Monthly earnings rate = $3575

Monthly hours worked = 35 Hrs

Hourly rate = 3575 / 35 = $102.14 per hr

Overtime rate between 35 and 40 hrs = $102.14

Overtime rate beyond 40 hrs = 150% of $102.14

= 1.5 x 102.14

= $153.21

For the given scenario,

A) The regular semi-monthly earnings = $3575 / 2

= $1787.50

B) Overtime = 10 hrs

Before 40 hrs = 8 hrs

After 40 hrs = 2 hrs

The Overtime earnings = Overtime earnings before 40 hrs + Overtime earnings after 40 hrs

= Sum (number of hours x hourly overtime rate)

= (8 x 102.14) + (2 x 153.21)

= $1123.54

C) Total Earnings = The regular semi-monthly earnings + Overtime earnings

= 1787.50 + 1123.54

= $2911 (Approx)